# Feature Coding in the Era of Large Models: Dataset, Test Conditions, and Benchmark

### v1

@@ Abstract
Large models have achieved remarkable performance across various tasks, yet they incur significant computational costs and privacy concerns during both training and inference. Distributed deployment has emerged as a potential solution, but it necessitates the exchange of intermediate information between model segments, with feature representations serving as crucial information carriers. To optimize information exchange, feature coding methods are applied to reduce transmission and storage overhead. Despite its importance, feature coding for large models remains an under-explored area. In this paper, we draw attention to large model feature coding and make three contributions to this field. First, we introduce a comprehensive dataset encompassing diverse features generated by three representative types of large models. Second, we establish unified test conditions, enabling standardized evaluation pipelines and fair comparisons across future feature coding studies. Third, we introduce two baseline methods derived from widely used image coding techniques and benchmark their performance on the proposed dataset. These contributions aim to advance the field of feature coding, facilitating more efficient large model deployment. All source code and the dataset will be made available on GitHub.
Figure 1: Three application scenarios for feature coding in large model deployment. Cloud-centralized training: Features generated by foundation models are encoded and stored in a feature database. This database supports various downstream tasks, reducing computational demands by avoiding repeated inference of large foundation models. Cloud-edge distributed training: To protect user data privacy, raw data is first processed by an edge model to generate features, which are then encoded and transmitted to the cloud. The cloud decodes these features to complete the remaining forward propagation. Cloud-edge distributed inference: A large model is partitioned across multiple parts, with portions offloaded to edge devices to distribute the computational load. The transmitted features ensure privacy protection and enable customized downstream tasks. Please refer to Sec. 3 for detailed illustrations.
@@ 1 Introduction
Large models have revolutionized artificial intelligence (AI) with their impressive performance. However, they require substantial training data and extensive computational resources, posing significant challenges for practical deployment [18]. To address these challenges, distributed deployment has been proposed as a promising solution [32, 37, 43, 46, 22, 14]. By distributing the computational workload across multiple platforms, this approach reduces resource strain on individual systems and enables the integration of private client data into training while maintaining data security [43, 4]. As model scale continues to grow, distributed deployment is expected to become a mainstream strategy for large model applications.
@@
In distributed deployment, effective information exchange between model segments is essential. During forward propagation, features act as information carriers, passing features generated by early segments to later segments to complete inference. Given the vast amount of data in large model applications, encoding features to a smaller size to reduce storage and transmission burden is necessary. Analogous to image and video coding, feature coding has been proposed to address this problem [6, 17, 16, 38].
@@
Existing feature coding research has three key limitations: narrow model types, limited task diversity, and constrained data modalities. Most existing feature coding efforts focus on small-scale, convolution-based networks, such as ResNet, ignoring large models. Although discriminative models, such as DINOv2 [28] and SAM [20], represent a significant branch of large models, they are still underrepresented in feature coding research. Generative models also have received little attention in this area. Furthermore, existing research primarily targets visual features, with limited integration of features from large language models (LLMs) based on textual inputs. This gap indicates a need for broader exploration across both model and data modalities to advance feature coding.
@@
In this work, we shift feature coding research from small-scale to large-scale models. To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity. Using this dataset, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
@@
To establish a foundation for large model feature coding, we propose two baseline methods and evaluate them on our dataset. The baselines leverage two primary image coding techniques: the handcrafted VVC Intra coding [2] and the learned Hyperprior method [1]. For each method, we introduce pre-processing and post-processing modules tailored to feature coding. Our analysis provides insights into baseline performance and highlights promising research directions.
In summary, our main contributions are as follows:
@@
  • We emphasize the importance of feature coding in large model deployment, identifying relevant application scenarios and extending feature coding research from small- to large-scale models.
@@
  • We construct a diverse feature dataset encompassing three model types, five tasks, and two data modalities, along with unified test conditions to support future feature coding research.
  • We propose and benchmark two feature coding methods on the dataset, offering baseline evaluations and insights for future research in this field.
@@
DINOv2 1.1B Image Classification $40^{th}$ ViT Block $257\times 1536$ ImageNet, 100 samples
Semantic Segmentation $40^{th}$ ViT Block $2\times 1370\times 1536$ VOC2012, 20 samples
Depth Estimation $10^{th}$, $20^{th}$, $30^{th}$, $40^{th}$ ViT Blocks $2\times 4\times 1611\times 1536$ NYU-Depth-v2, 16 samples
Llama3 8B Common Sense Reasoning $32^{th}$ Decoder Layer $N\times 4096$ Arc-Challenge, 100 samples
SD3 8B Text-to-Image Synthesis Input Layer of VAE Decoder $16\times 128\times 128$ COCO2017, 100 samples
@@ 2.1 Feature coding
Visual signal coding encodes and reconstructs the original visual data [33, 44, 15, 25, 34]. Unlike visual signal coding, feature coding refers to the process of encoding features into bitstreams. Originally introduced within the Coding for Machines framework [42, 10], feature coding is one of two primary branches alongside visual coding. One step in time ahead of feature coding, visual signal coding has already been explored in the context of large models [35].
Despite recent progress, feature coding research currently faces three limitations. First, there is a restricted variety of model types in use. The existing feature extraction relies on small-scale, CNN-based models such as ResNet series [23, 31, 19, 24, 3]. However, as Transformer architectures now dominate large model research [36, 11, 12, 28], it is crucial to investigate feature coding for Transformer-based models to capture the rich representations they provide. Second, the field is constrained by the limited scope of task types. Existing studies mainly address discriminative tasks like image classification and segmentation [7, 13, 45, 41, 5, 27], while generative tasks remain largely unexplored. Given the significance of generative models in modern AI, extending feature coding to include these tasks, such as image synthesis, is essential. The third limitation is the constrained range of source modalities. Current feature coding research predominantly focuses on features from visual data, overlooking the increasingly important features derived from textual data, particularly in light of advancements in LLMs. Expanding to multiple modalities is crucial as the diversity of AI applications grows.
Additionally, two core challenges hinder progress in feature coding for large models: the lack of a public test dataset and unified test conditions. Without a standardized dataset and evaluation conditions, fair comparisons across methods are not feasible. A public dataset ensures consistent input data across studies, while unified test conditions establish a consistent benchmark for evaluating performance.
These limitations motivate us to develop a diverse test dataset and formulate unified test conditions to advance feature coding research.
@@ 2.2 Large model deployment
User privacy presents another critical challenge in client services [4, 26, 40]. Split learning has been introduced as a solution, effectively protecting training data privacy by dividing computation between clients and clouds [43].
In distributed deployment, efficient information exchange between model segments is essential. However, the academic focus has largely overlooked the storage and transmission costs of exchanging these features. Motivated by this gap, our work encodes features into bitstreams to streamline storage and transmission, addressing an essential component of efficient distributed deployment for large models.
@@ 3 Feature coding in large models
This section presents three application scenarios of feature coding in large model deployment, illustrated in Fig. 1. Each scenario highlights the role of feature coding in optimizing deployment efficiency.
@@ 3.1 Large model training
Large models achieve their high performance through extensive training on large-scale datasets. Due to the limited computational capacity of edge devices, cloud-centralized training is commonly utilized. However, in sensitive domains like healthcare, high-quality public data may be scarce, necessitating access to private data through cloud-edge collaborative training.
@@ 3.1.1 Cloud-centralized training
In cloud-centralized training, source databases are maintained in the cloud, where intensive computational demands is the main challenge. Increasingly, complex tasks are addressed through the collaborative use of multiple models. For instance, in vision-language tasks, visual data is initially processed by a vision model, and its features are subsequently passed to an LLM for advanced reasoning. Similarly, foundation models often serve as core backbones for a range of tasks. For example, features extracted from DINOv2 are repurposed for downstream tasks like segmentation [9]. In such collaborative workflows, features generated by one model become input to other models. However, during training, each data sample may undergo repeated inference through the foundation model, leading to a substantial unnecessary computational load.
To mitigate this, an effective solution involves generating features once, storing them, and reusing them across tasks. The workflow is outlined on the left of Fig. 1. First, a foundation model extracts features from the source database, which are then encoded and stored in a feature database. For subsequent task-specific training, only the precomputed feature database is needed, enabling efficient access to features without redundant feature extraction. This significantly reduces computational load and accelerates the training of downstream tasks. However, raw feature data requires substantial storage space, especially with large datasets. To address this, we propose using feature coding to compress raw features into compact bitstreams, reducing storage demands and facilitating efficient, large-scale training.
@@ 3.1.2 Cloud-edge distributed training
In addition to computational resources, large model training demands access to large-scale datasets, which are often only available to a few giant high-tech companies. Furthermore, sensitive data, such as medical records and financial information, cannot be freely shared due to privacy concerns and regulatory constraints. To leverage distributed, privately owned data, split learning has emerged as a viable solution [37]. Split learning divides the training process into two phases, with the bulk of the computational load handled by the cloud, while edge devices, which are often resource-constrained, handle a smaller portion of the task. This approach minimizes the computational demand on the edge while enabling collaboration with powerful cloud resources.
The cloud-edge distribution training pipeline is depicted in the second diagram of Fig. 1. The whole pipeline is divided into two parts: edge models and cloud-side large models. During forward propagation, the edge model processes private data to generate features, which are then encoded into bitstreams and transmitted to the cloud. Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the large model. In the backward pass, gradients are computed and backpropagated along the reverse path. To optimize this process (reduce bandwidth usage and minimize transmission time), it is critical to apply feature coding before transmission.
@@ 3.2 Cloud-edge distributed inference
Once large models are trained, they are deployed in products to provide client services. For a product or service to succeed, two primary concerns must be addressed: supporting a high volume of simultaneous cloud server access and protecting user privacy. For the first concern, managing computational load is essential, particularly in generative applications that involve repetitive diffusion processes. A practical solution is to offload part of this computational demand to edge/client devices. For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec. 3.1.2, can apply.
The cloud-edge distributed inference pipeline is illustrated in the diagram on the right of Fig. 1. In the upload phase, source data is converted into features by an edge model on the client side. These features are then encoded and transmitted to the cloud, where they are further processed by large models. In the download phase, the processed features are sent back to the client to complete a specific task. Depending on the task head used, various functionalities can be performed with the returned features. For example, Stable Diffusion 3 (SD3) features derived from an image can be used to synthesize video or generate segmentation masks. This approach not only reduces the computational load on the cloud server but also enhances data privacy by keeping raw user data on the edge/client side.
@@
Figure 3: Visualization of the original feature blocks and their corresponding DCT blocks (absolute values). In the feature domain, the proposed features exhibit higher vertical redundancy, whereas the existing features vary smoothly in both directions. In the DCT domain, the proposed features display more dispersed energy distributions. These distinct redundancy characteristics indicate different coding properties.
@@ 4.1.1 Model and task selection
This study expands the scope of feature coding research by encompassing discriminative, generative, and hybrid models. While prior work has primarily concentrated on discriminative, convolution-based models, the advent of Transformer architectures and generative models calls for a more comprehensive dataset to capture a wider range of tasks and feature types. To meet this need, we select three representative models: DINOv2 [28], Llama3 [11], and SD3 [12]. Llama3 and SD3 represent two core categories within generative models, highlighting their respective strengths in different input-output modalities. Specifically, DINOv2 and Llama3 handle visual and textual inputs, while SD3 processes textual inputs to generate visual outputs. The selected models offer comprehensive feature representations across visual, textual, and cross-modal applications.
For DINOv2, we include the image classification, semantic segmentation, and depth estimation tasks. Llama3 addresses the common sense reasoning task, and SD3 focuses on text-to-image synthesis. This selection covers three visual tasks, one textual task, and one text-to-visual task, forming a versatile foundation for feature coding analysis across diverse task and model types.
@@ 4.1.2 Split point decision
This section describes the selection of split points for each model. For DINOv2, we define two types of split points: 1) $SP_{DS}$, the output of the $40^{th}$ ViT layer, and 2) $SP_{DM}$, which aggregates outputs from the $10^{th}$, $20^{th}$, $30^{th}$, and $40^{th}$ ViT layers. $SP_{DS}$ is used for tasks utilizing single-layer features, while $SP_{DM}$ supports tasks that require multi-layer features. In this study, $SP_{DS}$ is applied to image classification and semantic segmentation, and $SP_{DM}$ to depth estimation. The resulting features are designated as $F_{Cls}$, $F_{Seg}$, and $F_{Dpt}$, respectively.
For Llama3, the split point $SP_{G}$ is set at the output of the $32^{nd}$ decoder layer. Features extracted at $SP_{G}$ are versatile, supporting both task-specific heads and integration with other large models [39]. In the common sense reasoning task, $SP_{G}$ produces features of size $N\times 4096$, where $N$ represents the number of tokens.
For SD3, we establish the input layer of the VAE decoder as the split point $SP_{H}$. This split point generates features of size $16\times 128\times 128$, denoted as $F_{TTI}$, which are then passed to the VAE decoder for image synthesis.

### v3

@@ Abstract
Large models have achieved remarkable performance across various tasks, yet they incur significant computational costs and privacy concerns during both training and inference. Distributed deployment has emerged as a potential solution, but it necessitates the exchange of intermediate information between model segments, with feature representations serving as crucial information carriers. To optimize information exchange, feature coding is required to reduce transmission and storage overhead. Despite its importance, feature coding for large models remains an under-explored area. In this paper, we draw attention to large model feature coding and make three fundamental contributions. First, we introduce a comprehensive dataset encompassing diverse features generated by three representative types of large models. Second, we establish unified test conditions, enabling standardized evaluation pipelines and fair comparisons across future feature coding studies. Third, we introduce two baseline methods derived from widely used image coding techniques and benchmark their performance on the proposed dataset. These contributions aim to provide a foundation for future research and inspire broader engagement in this field. To support a long-term study, All source code and the dataset are made available at https://github.com/chansongoal/FCM-LM/tree/master.
Figure 1: Three application scenarios for feature coding in large model deployments. Cloud-centralized training: Features generated by foundation models are encoded and stored in a feature database. This database supports various downstream tasks, reducing computational demands by avoiding repeated inference of large foundation models. Cloud-edge distributed training: To protect user data privacy, raw data is first processed by an edge model to generate features, which are then encoded and transmitted to the cloud. The cloud decodes these features to complete the remaining forward propagation. Cloud-edge distributed inference: A large model is partitioned across multiple parts, with portions offloaded to edge devices to distribute the computational load. The transmitted features ensure privacy protection and enable customized downstream tasks. Please refer to Sec. 3 for detailed illustrations.
@@ 1 Introduction
Large models have revolutionized artificial intelligence (AI) with their impressive performance. However, they require substantial training data and extensive computational resources, posing significant challenges for practical deployments [19]. To address these challenges, distributed deployment has been proposed as a promising solution [35, 40, 46, 49, 23, 14]. By distributing the computational workload across multiple platforms, this approach reduces the resource strain on individual systems and enables the integration of private client data into training while maintaining data security [46, 4]. As model scale continues to grow, distributed deployment is expected to become a mainstream strategy for large model deployments.
@@
In distributed deployments, effective information exchange between model segments is essential. During forward propagation, features generated by early segments are passed to later segments to complete the inference. Given the vast amount of data in large model applications, it is necessary to encode features to a smaller size to reduce storage and transmission burden. Analogous to image and video coding, feature coding has been proposed to address this problem [6, 18, 16, 41].
@@
Existing feature coding research faces three key limitations: narrow model types, limited task diversity, and constrained data modalities. Most existing feature coding efforts focus on small-scale convolution-based networks, such as ResNet, ignoring large models. Although existing research emphasizes discriminative models, large models such as DINOv2 [31] and SAM [21] remain underrepresented. Generative models also have received little attention in this domain. Furthermore, existing research primarily targets visual features and neglects the investigation of textual features from large language models (LLMs). This gap indicates a need for broader exploration across both models and data modalities to advance feature coding.
@@
In this work, we shift the feature coding research from small-scale to large-scale models. To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity. Furthermore, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
@@
To lay the groundwork for large model feature coding, we propose two baseline methods and evaluate their applicability to the dataset. The baselines leverage two image coding methods: the handcrafted VVC Intra coding [2] and the learning-based Hyperprior method [1]. We introduce pre-processing and post-processing modules to adapt the baselines to feature characteristics. The resulting benchmark and analyses provide valuable insights into large model feature coding and identify promising directions for further exploration. In summary, our main contributions are as follows:
@@
  • We highlight the importance of feature coding in large model deployments, identify relevant application scenarios, and introduce a new research area: large model feature coding.
@@
  • We construct a comprehensive feature dataset encompassing three model types, five tasks, and two kinds of data modalities, along with unified test conditions to support long-term feature coding research.
  • We propose two baselines and evaluate their usability to the large model feature coding. We build a benchmark and the reported results suggest promising directions for further research.
@@
DINOv2 1.1B Image Classification (Cls) $40^{th}$ ViT Block $257\times 1536$ ImageNet, 500 samples
Semantic Segmentation (Seg) $40^{th}$ ViT Block $2\times 1370\times 1536$ VOC2012, 100 samples
Depth Estimation (Dpt) $10^{th}$, $20^{th}$, $30^{th}$, $40^{th}$ ViT Blocks $2\times 4\times 1611\times 1536$ NYU-Depth-v2, 80 samples
Llama3 8B Common Sense Reasoning (CSR) $32^{th}$ Decoder Layer $N\times 4096$ Arc-Challenge, 500 samples
SD3 8B Text-to-Image Synthesis (TTI) Input Layer of VAE Decoder $16\times 128\times 128$ COCO2017, 500 samples
@@ 2.1 Feature coding
Originally introduced within the Coding for Machines framework [45, 10], feature coding is one of two primary branches alongside visual coding. Unlike visual signal coding [36, 47, 15, 26, 37, 29], which encodes and reconstructs the original visual data, feature coding involves encoding features extracted from source data (e.g., images) into bitstreams. While visual coding has already been explored in the context of large models [38], feature coding remains underdeveloped in this domain.
Despite recent progress, feature coding research currently faces three key limitations. First, there is a restricted variety of model types in use. The existing feature extraction relies on small-scale convolution-based models such as ResNet series [24, 34, 20, 25, 3, 28, 17]. However, as Transformer architectures now dominate large model research [39, 11, 12, 31], it is crucial to investigate feature coding for Transformer-based models to capture the rich representations they provide. Second, the field is constrained by the limited scope of task types. Existing studies mainly address discriminative tasks like image classification and segmentation [7, 13, 48, 44, 5, 30], while generative tasks remain largely unexplored. Given the significance of generative models in modern AI, extending feature coding to include these tasks, such as image synthesis, is essential. The third limitation is the constrained range of source modality. Current feature coding research predominantly focuses on features extracted from visual data, overlooking the increasingly important features generated from textual data, particularly in light of advancements in LLMs. Expanding to multiple modalities is crucial as the diversity of AI applications grows.
Additionally, two core challenges hinder progress in feature coding for large models: the lack of a public test dataset and unified test conditions. Without a standardized dataset and evaluation conditions, fair comparisons across methods can not be conducted. A public dataset ensures consistent input data across studies, while unified test conditions establish a standard pipeline to evaluate performance.
These limitations motivate us to construct a diverse test dataset and formulate unified test conditions to advance feature coding research.
@@ 2.2 Large model deployment
User privacy protection presents another challenge in client services [4, 27, 43]. Split learning has been proposed as an effective solution to protect data privacy by partitioning models between clients and the cloud [46].
In distributed deployments, efficient information exchange between model segments is crucial. However, academic research has largely overlooked the storage and transmission costs associated with feature exchange. To address this gap, we propose to encode features into compact bitstreams to improve the efficiency of large model deployments in distributed environments.
@@ 3 Feature coding in large models
In Fig. 1, we illustrate three application scenarios of feature coding in large model deployments, highlighting their roles in optimizing the deployment efficiency of large models.
@@ 3.1 Large model training
Large models achieve their high performance through extensive training on large-scale datasets. Cloud-centralized training is commonly employed to leverage the vast computational resources and data in the cloud. However, in sensitive domains like healthcare, high-quality public data may be scarce, necessitating access to private data through cloud-edge collaborative training.
@@ 3.1.1 Cloud-centralized training
In cloud-centralized training, source databases are maintained in the cloud, where intensive computational demands become the main challenge. Complex tasks are increasingly addressed through the collaborative use of multiple models. For instance, in vision-language tasks, visual data is first processed by a vision model, and its features are then passed to an LLM for advanced reasoning. Similarly, foundation models often serve as core backbones for various tasks. For example, features extracted from DINOv2 are repurposed for downstream tasks like segmentation [9]. In such collaborative workflows, features generated by one model become inputs for others. However, during training, each data sample may undergo repeated inference, resulting in substantial and unnecessary computational costs.
To address this, an effective solution is to generate features once, store them, and reuse them across tasks. The workflow is outlined on the left of Fig. 1. First, a foundation model extracts features from the source database, which are then encoded and stored in a feature database. For subsequent task-specific training, only the precomputed feature database is needed, enabling efficient access to features without redundant feature extraction. This approach significantly reduces computational load and accelerates downstream task training. However, raw feature data consumes considerable storage space, particularly with large datasets. Therefore, encoding raw features into compact bitstreams becomes essential to reduce storage demands and support efficient large model training.
@@ 3.1.2 Cloud-edge distributed training
In addition to computational resources, large model training requires access to large-scale datasets, which are often only available to a few major high-tech companies. Furthermore, sensitive data, such as medical records and financial information, cannot be freely shared due to privacy concerns and regulatory constraints. To leverage distributed, privately owned data, split learning has emerged as a viable solution [40]. Split learning divides the training process into two phases, with the bulk of the computational load handled by the cloud, while edge devices, which are often resource-constrained, handle a smaller portion of the task. This approach minimizes the computational demand on the edge while enabling collaboration with powerful cloud resources.
The cloud-edge distribution training pipeline is depicted in the middle of Fig. 1. The whole pipeline is divided into two parts: edge models and cloud models. During forward propagation, the edge model processes private data to generate features, which are then encoded into bitstreams and transmitted to the cloud. Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the cloud model. In the backward pass, gradients are computed and backpropagated along the reverse path. To optimize this process (reduce bandwidth costs and minimize latency), it is critical to apply feature coding before transmission.
@@ 3.2 Cloud-edge distributed inference
Once large models are trained, they are deployed in products to provide client services. For a product or service to succeed, two primary concerns must be addressed: supporting a high volume of simultaneous cloud server access and protecting user privacy. For the first concern, managing computational load is essential, particularly in generative applications that involve repetitive diffusion processes. A practical solution is to offload part of this computational demand to edge devices. For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec. 3.1.2, can apply.
The cloud-edge distributed inference pipeline is illustrated on the right of Fig. 1. In the upload phase, source data is converted into features by an edge model on the client side. These features are then encoded and transmitted to the cloud, where they are further processed by large models. In the download phase, the processed features are sent back to the client to complete a specific task. Depending on the task head used, various functionalities can be performed with the returned features. For example, Stable Diffusion 3 (SD3) features derived from an image can be used to synthesize video or generate segmentation masks. This approach not only reduces the computational load on the cloud server but also enhances data privacy by keeping raw user data on the edge side.
@@
Figure 3: Visualization of the original feature blocks and their corresponding DCT blocks (absolute values). In the feature domain, the proposed features exhibit higher vertical redundancy, whereas the existing features vary smoothly in both directions. In the DCT domain, the proposed features display more dispersed energy distributions. These distinct redundancy characteristics indicate different coding properties. We use $7\times 7$ blocks for $C5$ and $C5$-ReLU and $16\times 16$ blocks for other features.
@@ 4.1.1 Model and task selection
Given the vast number of large models available, it is impractical to include all of them in a single research paper. Therefore, we focus on selecting representative models. To ensure the dataset’s representativeness and long-term relevance, we choose one widely recognized model for each type of large model. Specifically, we select DINOv2 [31], Llama3 [11], and SD3 [12] as representatives of discriminative models, generative models, and hybrid models, respectively. The trend of adopting Transformer architectures in large models is expected to persist in the coming years. As a result, the selected models are anticipated to retain their research value and practical utility in the near future. DINOv2 and Llama3 handle visual and textual inputs, respectively, while SD3 processes textual inputs to generate visual outputs. Together, these models provide comprehensive feature representations across visual, textual, and cross-modal applications.
For DINOv2, we include image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt) tasks. Llama3 is applied to the common sense reasoning task (CSR), and SD3 is used for text-to-image synthesis (TTI). The selection covers three visual tasks, one textual task, and one text-to-visual task, forming a comprehensive foundation for analyzing feature coding across diverse tasks and model types.
@@ 4.1.2 Split point decision
The split points are decided to support various downstream tasks. For DINOv2, we define two types of split points: 1) $SP_{DS}$, the output of the $40^{th}$ ViT layer, and 2) $SP_{DM}$, which aggregates outputs from the $10^{th}$, $20^{th}$, $30^{th}$, and $40^{th}$ ViT layers. $SP_{DS}$ is used for tasks utilizing single-layer features, while $SP_{DM}$ supports tasks that require multi-layer features. In this study, $SP_{DS}$ is applied to Cls and Seg, and $SP_{DM}$ to Dpt. The corresponding features are designated as $F_{Cls}$, $F_{Seg}$, and $F_{Dpt}$, respectively.
For Llama3, the split point $SP_{G}$ is set at the output of the $32^{nd}$ decoder layer. Features extracted at $SP_{G}$ can support both task-specific heads and integration with other large models [42]. In CSR, $SP_{G}$ produces $N\times 4096$ features, where $N$ is the number of tokens.
For SD3, we take the input layer of the VAE decoder as the split point $SP_{H}$. This split point generates features of size $16\times 128\times 128$, denoted as $F_{TTI}$, which are then passed to the VAE decoder for image synthesis.